### v1

# Birkhoff’s Theorem and Uniqueness: A Peak Beyond General Relativity

Abstract
In General Relativity, Birkhoff’s theorem asserts that any spherically symmetric vacuum solution must be static and asymptotically flat. In this paper, we study the validity of Birkhoff’s theorem for a broad class of modified gravity theories in four spacetime dimensions, including quadratic and higher-order gravity models. We demonstrate that the Schwarzschild spacetime remains the unique Einstein branch solution outside any spherically symmetric configuration of these theories. Consequently, unlike black holes, the breakdown of junction conditions at the surface of the star further implies that the actual spacetime metric outside a horizonless star in these modified theories cannot simultaneously be spherically symmetric and remain within the Einstein branch. This insight offers a unique observational probe for theories beyond General Relativity.
I Introduction
Even after a century of its inception, Einstein’s theory of General Relativity (GR) remains the most successful classical framework for describing gravitational phenomena across a vast range of length scales. The predictions of GR have been extensively validated using both classical tests and modern observations — spanning over solar system tests to the unprecedented precision offered by cosmological surveys Will (2014). Lately, these tests involve the state-of-the-art gravitational wave detections by the LIGO-Virgo-KAGRA (LVK) collaboration Abbott et al. (2016a, b, c, 2023, 2021); Cahillane and Mansell (2022) and shadow observations by the Event Horizon Telescope (EHT) collaboration Akiyama et al. (2019); Event Horizon Telescope Collaboration (2022). In light of these remarkable developments, significant efforts have been put to investigate possible deviations from GR solutions, particularly from the Schwarzschild paradigm, which has been widely used in both classical and modern tests of GR to model spacetime outside spherically symmetric objects when rotational effects are negligible Will (2014). These investigations are critical as they probe the limits of GR and explore the potential for new physics in extreme gravity environments where deviations from standard GR solutions might become detectable.
The wide-ranging applications of the Schwarzschild metric are primarily supported by two important facts. Firstly, understanding a spherically symmetric solution like Schwarzschild provides valuable input for finding more realistic rotating solutions. Indeed, there exist well-defined techniques, such as the Newman-Janis algorithm Newman and Janis (1965), to construct rotating solutions from spherically symmetric ones. Secondly, one of the most important results in GR, namely Birkhoff’s theorem, establishes the uniqueness of Schwarzschild solution outside a spherically symmetric object Jebsen (1921); Birkhoff and Langer (1923). In addition to its foundational role, Birkhoff’s theorem has profound observational consequences. For example, along with the regularity of the spacetime at the centre, it implies that the metric inside an empty spherical cavity must be the Minkowski. Moreover, Birkhoff’s theorem claims the absence of time-dependent vacuum solutions outside a star due to its spherical pulsations. Hence, such pulsations cannot alter the exterior metric from that of Schwarzschild, thereby producing no gravitational radiation.
Over the years, several generalizations of this novel theorem have extended its domain of applicability. The usual proof of Birkhoff’s theorem in GR requires that the metric be at least twice continuously differentiable ($C^{2}$) for calculating the Ricci tensor. However, the theorem is also applicable under a weaker assumption of continuous and piecewise differentiability of the metric Bergmann et al. (1965). The validity of the theorem has also been extended to include electromagnetic fields, for which Reissner-Nordström is the unique spherically symmetric solution of the Einstein-Maxwell equations D’Inverno (1992). More generalizations of Birkhoff’s theorem involve the inclusion of cosmological constant Schleich and Witt (2010), as well as extensions to lower and higher spacetime dimensions Kiem (1994); Schmidt (1997); Bronnikov and Melnikov (1995); Ayon-Beato et al. (2004); Keresztes and Gergely (2008) and mild perturbations away from exact spherical symmetry and vacuum Goswami and Ellis (2011, 2012).
Since Birkhoff’s theorem is a direct consequence of GR field equations, its applicability in the modified gravity landscape is not generally guaranteed. However, the importance of carefully assessing its validity in beyond-GR theories can hardly be exaggerated. In fact, there are strong theoretical and observational reasons to believe that GR might receive higher curvature corrections Penrose (1965); Donoghue (1994); Wald (1997); Peebles and Ratra (2003); Bertone et al. (2005); Clifton et al. (2012); Barack et al. (2019), which becomes particularly important in strong-field regimes Berti et al. (2015); Berti (2019). Thus, it is both interesting and important to study the status of various novel results of GR, like Birkhoff’s theorem, within the framework of these modified theories. Such studies can provide insights into potential universal features of the beyond-GR landscape. Motivated by these considerations, Birkhoff’s theorem has already been extended to Lovelock gravity Zegers (2005); Deser and Franklin (2005), and more generally quasi-topological gravity Oliva and Ray (2011); Cisterna et al. (2017); Oliva and Ray (2010); Bueno et al. (2024) in $D>4$ dimensions. Additionally, Birkhoff’s theorem has been explored in the context of scalar-tensor theories Faraoni (2010), Einstein-æther theory Chan et al. (2024) and Horava-Lifshitz gravity Devecioglu and Park (2019). It is also known to hold in $f(R)$ gravity under certain conditions Capozziello and Saez-Gomez (2012); Ravindranath et al. (2018); Nzioki et al. (2014). However, a comprehensive approach to extend its validity to other modified theories is still lacking.
In this work, we aim to bridge this crucial gap for the first time by establishing Birkhoff’s theorem for a wide class of beyond-GR theories in $4$-dimensions (4D). The Lagrangians of these theories involve both Ricci scalar and Ricci tensor invariants, which will be discussed in more detail in the subsequent sections. Unlike in GR, such modified theories lead to multiple branches of solutions, which may be either analytic or non-analytic functions of higher curvature coupling constants. However, the remarkable observational agreement with GR strongly suggests that only the Einstein branch is astrophysically relevant. The solutions in this branch are smoothly connected to GR in the limit of vanishing coupling constants. In other words, all geometrical quantities (such as metric components and curvatures) can be written as a linear combination of analytical terms in couplings. Furthermore, stationary BHs in the Einstein branch are promising thermodynamic candidates beyond GR, as they have been shown to satisfy the zeroth law Ghosh and Sarkar (2020); Bhattacharyya et al. (2022).
These novel properties motivate us to explore the extension of Birkhoff’s theorem within the Einstein branch of a class of modified theories discussed below. In particular, we shall show that Schwarzschild solution remains the unique vacuum solution outside spherically symmetric objects in the Einstein branch of these theories. Similar to the case of GR, the staticity and asymptotic flatness of the exterior spacetime follow from the spherical symmetry. We shall also discuss several important implications of this result, which are of utmost observational relevance. For instance, we shall argue that as a direct consequence of this powerful theorem, the actual spacetime metric outside a horizonless star in the theories under consideration cannot be both spherically symmetric and remain within the Einstein branch due to the failure of junction conditions at the surface of the star Khakshournia and Mansouri (2023). In other words, unlike GR, the spacetime outside a spherically symmetric star in these theories must deviate from the Einstein branch. This deviation, in turn, provides a novel observational probe into the nature of beyond-GR theories.
Throughout this work, we shall set $G=c=1$. We begin our analysis with quadratic gravity (QG) in Sect.II as a warm-up, which we then generalize to more comprehensive beyond-QG theories in Sect. III. Finally, we conclude in Sect. IV by outlining potential future generalizations of our findings.
II Birkhoff’s theorem in Quadratic Gravity
According to the Gauss-Bonnet theorem, the most general gravity theory in $4$-dimensions including terms up to quadratic order in curvatures is the so-called quadratic gravity  Stelle (1977, 1978) with the Lagrangian density: $\mathcal{L}=R+\alpha\,R^{2}+\beta\,R_{\mu\nu}R^{\mu\nu}$. While QG is traditionally considered problematic due to the presence of massive spin-2 ghost modes (suggesting perturbative instability), it can still be treated as a “healthy” classical theory of gravity. In fact, some recent works have demonstrated that QG is free from causality issues Edelstein et al. (2021), positive energy theorem holds (implying the nonlinear stability), and possesses well-defined dynamical properties Held and Lim (2023).
It is crucial to recognize that QG, being a higher-curvature theory, may have multiple branches of solutions Holdom (2002); Daas et al. (2023); Lü et al. (2015); Lu et al. (2015); Lü et al. (2015); Kokkotas et al. (2017); Pravda et al. (2017); Lü et al. (2017); Podolský et al. (2020). However, as emphasized in the Introduction section, the remarkable observational consistency of GR dictates that only the Einstein branch is astrophysically relevant. The solutions in this branch are analytically connected to GR in the limit of $\{\alpha,\,\beta\}\to 0$. Consequently, all geometrical quantities can be expressed as a combination of terms proportional to $\alpha^{n_{1}}\beta^{n_{2}}$, with $n_{i}$’s being non-negative integers for all values of the couplings. One must note the subtle difference between the above claim and the Taylor expansion (which is not what we are doing) about $\{\alpha,\,\beta\}=0$.
The QG field equations in vacuum are given by $E_{\mu\nu}:=G_{\mu\nu}+\alpha\,H_{\mu\nu}+\beta\,I_{\mu\nu}=0$, where $G_{\mu\nu}$ is the Einstein tensor, and the explicit expressions of $\{H_{\mu\nu},\,I_{\mu\nu}\}$ can be found in Ref. Stelle (1977). The most general spherically symmetric solution (characterizing spacetime outside spherical objects) of this theory in $4$-dimensions can be written as follows:
$ds^{2}=-e^{f(r,t)}\,dt^{2}+e^{g(r,t)}\,dr^{2}+r^{2}\,d\Omega^{2}_{(2)}\,.$ (1)
Note that the cross terms such as $g_{t\theta}$, $g_{t\phi}$, $g_{r\theta}$, and $g_{r\phi}$ are zero because of spherical symmetry. Furthermore, the metric can always be diagonalized in the $t-r$ sector to eliminate the $g_{tr}$ term. Since we are only interested in solutions in the Einstein branch, one must have
$f(t,r)=\mathrm{ln}\Big{(}1-\frac{2M}{r}\Big{)}+\epsilon\,f^{(1)}(t,r)+\epsilon^{2}\,f^{(2)}(t,r)+\cdots\,,$ (2)
$g(t,r)=-\mathrm{ln}\Big{(}1-\frac{2M}{r}\Big{)}+\epsilon\,g^{(1)}(t,r)+\epsilon^{2}\,g^{(2)}(t,r)+\cdots\,.$ (3)
Here, $\epsilon$ is a book-keeping parameter (at the end we shall set it to $1$) tracking the order of $\{\alpha^{n},\,\beta^{n}\}$ in the above expressions. More explicitly, the term proportional to $\epsilon^{n}$ contains all combinations of $\alpha^{n_{1}}\,\beta^{n_{2}}$ such that $n_{1}+n_{2}=n$ ($n_{i}\geq 0$). We have taken the $\epsilon^{0}$-th order solution to be same as GR as Birkhoff’s theorem holds in vacuum GR. Thus, to prove the Birkhoff’s theorem in QG, we must first demonstrate that the metric Eq. (1) is actually static. For this purpose, let us now study the properties of various components of the field equations order-by-order in $\epsilon$ and as functions of various metric coefficients in the Einstein branch.
(a) Till $\epsilon^{1}$ order: In this order, it is trivial to check that $E_{tr}=(\epsilon/r)\,\partial_{t}g^{(1)}$. Setting this to zero immediately implies that $g^{(1)}(t,r)$ must be independent of $t$. Then, substituting this in the expression of $\partial_{t}E_{rr}\propto\partial_{t}\partial_{r}f^{(1)}(t,r)=0$ leads to the condition that $f^{(1)}(t,r)=h^{(1)}(t)+j^{(1)}(r)$. Then, one can absorb the $h(t)$ part into a redefined time coordinate. This completes our proof at $\epsilon^{1}$ order.
(b) Till $\epsilon^{2}$ order: Similarly, one can check that $E_{tr}=(\epsilon^{2}/r)\,\partial_{t}g^{(2)}(t,r)=0$. It implies that $g^{(2)}(t,r)$ is independent of $t$. Substituting this back in the expression of $\partial_{t}E_{rr}\propto\partial_{t}\partial_{r}f^{(2)}(t,r)=0$, we get $f^{(2)}(t,r)=h^{(2)}(t)+j^{(2)}(r)$. Finally, we can absorb $h(t)$ into a redefined time coordinate.
In fact, one can inductively show that the above calculation follows at all orders in $\epsilon^{n}$, finishing our proof that the metric in Eq.(1) must be static. Then, the question remains that what are these metric coefficients that solve QG field equations? To answer this question, let us again consider the $tt$ and $rr$ components of the QG field equation. Up to quadratic order in $\epsilon$, we get the following expressions:
$g_{tt}(r)\approx-\Big{(}1-\frac{2M}{r}\Big{)}-\epsilon\,\frac{g_{c1}-2\,f_{c1}\,M+f_{c1}\,r}{r}\,-\frac{\epsilon^{2}}{r}\times$
$\left[f_{c1}\,g_{c1}-M\,(f_{c1}^{2}+2\,f_{c2})+f_{c1}^{2}\,r/2\,+(g_{c2}+f_{c2}\,r)\right]\,,$ (4)
$g_{rr}(r)\approx\Big{(}1-\frac{2M}{r}\Big{)}^{-1}-\epsilon\,\frac{g_{c1}\,r}{(r-2M)^{2}}+\frac{\epsilon^{2}}{(r-2M)^{3}}\times$
$\kern 100.00015pt\left[r(g_{c1}^{2}+2\,g_{c2}\,M-g_{c2}\,r)\right]\,.$ (5)
Here, $g_{c\,n}$ and $f_{c\,n}$ are some constants depending on $\alpha^{n_{1}}\beta^{n_{2}}$, such that $n_{1}+n_{2}=n$ ($n_{i}\geq 1$). One can also check that, till quadratic order, $g_{tt}\,g_{rr}\approx-1-f_{c1}\,\epsilon-(f_{c1}^{2}/2+f_{c2})\,\epsilon^{2}$. Therefore, by a redefinition of the time coordinate (absorbing some constants in $dt^{2}$), we can make $g_{tt}\,g_{rr}=-1$. In fact, going into this new time coordinate is equivalent to setting $f_{c\,n}=0$ for all $n\geq 1$. With this rearrangement, the metric components become
$-g_{tt}(r)=\Big{(}1-\frac{2M^{\prime}}{r}\Big{)}=g_{rr}^{-1}(r)\,,$ (6)
where $M^{\prime}=M-0.5(g_{c1}\,\epsilon+g_{c2}\,\epsilon^{2}+\cdots)$ is the redefined ADM mass. Hence, we have shown that in the Einstein branch, the unique spherically symmetric solution of QG is Schwarzschild. This completes our proof of Birkhoff’s theorem in QG. However, we shall see in Sec. IV that as a consequence of this result, when complied with the failure of junction conditions at the surface of a horizonless star Khakshournia and Mansouri (2023), the actual spacetime metric outside the star in QG cannot be simultaneously spherically symmetric and remain within the Einstein branch.
A few additional comments on this result are in order. It is well known that Schwarzschild metric is a solution of QG. However, what we have demonstrated is that Schwarzschild is the unique spherically symmetric vacuum solution (with staticity arising automatically) within the Einstein branch of QG. Notably, there exists a known spherically symmetric, Ricci-flat solution of QG, which is not analytic as it contains $\beta^{-1}$ terms and hence, it is not in the Einstein branch Pravda et al. (2017); Lü et al. (2017); Podolský et al. (2020). Furthermore, taking the trace of the QG field equation, we get $2(3\alpha+\beta)\Box R=R$. Here, the combination $2(3\alpha+\beta)$ is the inverse-squared mass of the massive scalar (or breathing) mode present in QG. However, since the solution we have found is uniquely Schwarzschild, having $R=0$, the breathing mode is non-propagating. This is consistent with expectations, as Birkhoff’s theorem implies that spherically symmetric pulsations cannot alter the metric outside the central spherical object.
III Extension beyond Quadratic Gravity
To generalize Birkhoff’s theorem beyond QG, we revisit our previous analysis from a different perspective. We consider a class of modified theories with vacuum field equations: $G_{\mu\nu}+\epsilon\,K_{\mu\nu}=0$, with the assumption that $K_{\mu\nu}$ is a functional of Ricci tensor and its covariant derivatives only, such that $K_{\mu\nu}[g^{Sch}_{\mu\nu}]=0$. However, this apriori does not guarantee that Schwarzschild is the unique spherically symmetric solution of these theories (which encompass both QG and $f(R)$-gravity) and staticity follows from the above field equations and spherical symmetry. However, it is exactly what we show now using a unified technique.
To establish the uniqueness of the Schwarzschild solution as the spherically symmetric vacuum solution of this class of modified gravity, we again consider the Einstein branch and work with the same metric ansatz as given in Eq.(1) along with Eq.(2). In the first order in $\epsilon$, we symbolically write the metric as $g_{\mu\nu}(t,r)=g_{\mu\nu}^{Sch}(r)+\epsilon\,g_{\mu\nu}^{(1)}(t,r)$. Substituting it in the field equations and using the fact that $K_{\mu\nu}[g_{\mu\nu}^{Sch}]=0$, we get $G_{\mu\nu}[g_{\mu\nu}^{Sch}+\epsilon\,g_{\mu\nu}^{(1)}]=\mathcal{O}(\epsilon^{2})$  [^1] This is distinct from the Taylor expansion about $\epsilon=0$., which is just GR field equations with the deviated metric. However, since Birkhoff’s theorem holds in GR, we must have $g_{\mu\nu}(t,r)=g_{\mu\nu}^{Sch}(r)$ till the first order in $\epsilon$, although the ADM mass will receive $\epsilon$-order correction as shown in the previous section. Then, $g_{\mu\nu}(t,r)=g_{\mu\nu}^{Sch}(r)+\epsilon^{2}\,g_{\mu\nu}^{(2)}(t,r)$ up to second order. The above argument can be easily extended to all orders in $\epsilon$, completing the proof. Therefore, if a modified theory of gravity supports Schwarzschild metric as a vacuum solution, then Birkhoff’s theorem holds and Schwarzschild remains the unique spherically symmetric vacuum solution in the Einstein branch (but see the caveat in Sec. IV when complied with the junction condition for a star).
A few important comments are in order. As a special case of the above construction, it is easy to see that GR vacuum shock wave (i.e., boosted BH solution to velocities tending to light’s speed producing a singular localized stress-energy tensor) solution is the unique Einstein branch shock in this class of modified gravity, extending Horowitz’s result beyond $\sigma$-model Horowitz and Steif (1990). Moreover, it follows that if $K_{\mu\nu}[g_{\mu\nu}^{Kerr}]=0$, then Kerr is the unique stationary and asymptotically flat vacuum BH solution in the Einstein branch of this particular class of modified gravity. Note that, unlike the spherically symmetric case, the stationarity and asymptotic flatness of Kerr does not follow from the field equations and need to be assumed.
IV Conclusion and Discussions
To date, the extension of Birkhoff’s theorem beyond GR remains limited. However, the theoretical importance and observational consequences of such an extension to other theories can be hardly overstated. This work aims to bridge this gap for the first time by establishing Birkhoff’s theorem for a broad class of beyond-GR theories, including QG. Our proof is valid in the Einstein branches of these theories, which is smoothly connected to GR in the limit of vanishing coupling constants.
Consequently, we demonstrate that the Schwarzschild solution remains the unique vacuum solution outside spherically symmetric objects in the Einstein branch of these theories. As in GR, the staticity and asymptotic flatness of the exterior spacetime follow as a direct consequence of spherical symmetry. This result has some profound implications. For instance, since the Schwarzschild metric has a vanishing Ricci scalar and tensor, it cannot be matched with any valid internal solution in the case of a horizonless object. Hence, the failure of junction conditions, which is now modified for beyond-GR theories Khakshournia and Mansouri (2023), suggest that the outside spacetime of a horizonless star cannot both be spherically symmetric and lie simultaneously in the Einstein branch.
Possible future extensions of our work include relaxing the assumption of asymptotic flatness and studying the validity of Birkhoff’s theorem for asymptotically de Sitter or ant-de Sitter-type spacetimes. To this end, we note that our argument needs modification for incorporating cases of non-vacuum solutions and those modified theories that imply $K_{\mu\nu}[g_{\mu\nu}^{Sch}]\neq 0$. Moreover, it will be interesting to study this theorem beyond the Einstein branches of various modified gravity. We leave these novel studies for our future endeavors.
V Acknowledgment
We thank Timothy Clifton for his valuable comments on an earlier version of this draft. We also appreciate the insightful discussions with Sumanta Chakraborty during the project’s initial stage. RG extends his gratitude to Sumanta Chakraborty for the invitation to the Indian Association for the Cultivation of Science (IACS) and to all members of the IACS Physics Department for their warm hospitality during his visit. The work of AC is supported by the National Postdoctoral Fellowship of SERB, ANRF, Government of India (File No.: PDF/2023/000550).